what does r equal in this equation : 8r+3=5r+12?
also what does p equal in this equation : 9p-14=4p+11?

Answer by arunpaul(104)

1st problem
8r + 3 = 5r + 12
or 8r - 5r = 12 - 3
or 3r = 9
or r = 9/3 = 3 (ans)
2nd problem
9p - 14 = 4p + 11
or 9p - 4p = 11 + 14
or 5p = 25
or p = 25/5 = 5 (Ans)